a. Obtain the linear trend equation for the following data on new checking accounts at Fair Savings Bank and use it to predict expected new checking accounts for periods 16 through 19. (Round your intermediate calculations and final answers to 2 decimal places.)

Period New Accounts Period New Accounts Period New Accounts
1 200 6 239 11 281
2 212 7 243 12 275
3 211 8 250 13 287
4 222 9 251 14 288
5 235 10 267 15 316
Y = + t
Y16 =      
Y17 =
Y18 =
Y19 =


b.Use trend-adjusted smoothing with α = .2 and β = .1 to smooth the new account data in part a. What is the forecast for period 16? Compute the initial trend estimate (Tt) for Period 5 as follows: (Period 4 data – Period 1 data) / 3. Then compute the initial trend-adjusted forecast (TAFt) for Period 5 as follows: Period 4 data + Initial trend estimate for Period 5. Then compute all remaining values (including the St value for Period 5) using the textbook formulas or Excel template. (Round the "Trend"values (Tt) to 3 decimal places and all other intermediate forecast values (TAFt and St) to 2 decimal places. Round your final answer to 2 decimal places.)


Forecast for period 16 =___  
Answer #1

a)

Period(x)

New Accounts(y)

xy

x2

1

200

200

1

2

212

424

4

3

211

633

9

4

222

888

16

5

235

1175

25

6

239

1434

36

7

243

1701

49

8

250

2000

64

9

251

2259

81

10

267

2670

100

11

281

3091

121

12

275

3300

144

13

287

3731

169

14

288

4032

196

15

316

4740

225

Total

120

3777

32278

1240

x-bar = Sum(x)/n = 120/15 = 8
y-bar = Sum(y)/n = 3777/15 = 251.8

b = (Sum(xy) – n*x-bar*y-bar)/(Sum(x2) – n*x-bar*x-bar)
= (32278-15*8*251.8)/(1240–15*8*8)
= 2062/280 = 7.364285714 = 7.36 (Rounded to 2 decimal place)

a = y-bar –b*x-bar
= 251.8 - 7.36*8 = 192.92

Regression equation is y = a + bx,
we get following linear trend equation by substituting values of a and b
Regression equation is y =192.92+7.36x

Y16 = 192.92+7.36*16 = 310.68

Y17 = 192.92+7.36*17 = 318.04

Y18 = 192.92+7.36*18 = 325.4

Y19 = 192.92+7.36*19 = 332.76

b)

Period(t)

New Accounts(A(t))

S(t)

T(t)

TAF(t)

1

200

2

212

3

211

4

222

5

235

222

7.333

229.33

6

239

230.46

7.446

237.91

7

243

238.13

7.468

245.60

8

250

245.08

7.416

252.50

9

251

252

7.366

259.37

10

267

257.70

7.199

264.90

11

281

265.32

7.241

272.56

12

275

274.25

7.410

281.66

13

287

280.33

7.277

287.61

14

288

287.49

7.265

294.76

15

316

293.41

7.130

300.54

303.63

7.439

311.07

Using trend adjusted exponential smoothing, F(t) = TAF(t-1) + alpha*(A(t-1) - TAF(t-1))

T(t) = T(t-1) + beta*(F(t) - TAF(t-1))

and TAF(t) = F(t) + T(t)

Forecast for period 16 = 311.07